### v1

# A patchy-particle 3-dimensional octagonal quasicrystal

Abstract
We devise an ideal 3-dimensional octagonal quasicrystal that is based upon the 2-dimensional Ammann-Beenker tiling and that is potentially suitable for realization with patchy particles. Based on an analysis of its local environments we design a binary system of 8- and 5-patch particles that in simulations assembles into a 3-dimensional octagonal quasicrystal. The local structure is subtly different from the original ideal quasicrystal possessing a narrower coordination-number distribution; in fact, the 8-patch particles are not needed and a one-component system of the 5-patch particles assembles into an essentially identical octagonal quasicrystal. We also consider a one-component system of the 8-patch particles; this assembles into a cluster with a number of crystalline domains, but which, because of the coherent boundaries between the crystallites, has approximate eight-fold order. We envisage that these systems could be realized using DNA origami or protein design.
Quasicrystals (QCs) have long range order, as exemplified by the sharp peaks in their diffraction patterns, but without a periodically repeating unit cell. QCs can thus have symmetries that are not possible for periodic crystals. The first example was discovered by Shechtman for a Al/Mn alloy and had icosahedral symmetry [1]. This discovery stimulated a search for further examples and quasicrystalline alloys with decagonal [2], dodecagonal [3, 4] and octagonal [5, 6] symmetry were quickly discovered. More recently, QCs have also been experimentally discovered in soft matter systems [7, 8, 9], but these have been limited to dodecagonal symmetry. Thus, discovering new ways to form QCs and increasing the possible repertoires of structures and symmetries is of significant interest.
Theory and simulations have the potential to significantly contribute to this goal by identifying how interparticle interactions can be designed to achieve quasicrystallinity. QCs involve multiple length scales that are related by a specific irrational ratio, e.g. the golden ratio in the case of icosahedral QCs. This feature is exploited in approaches that use isotropic potentials with multiple length scales [10, 11]. For example, a theoretical framework for designing such ultra-soft potentials has been developed [12, 13, 14]. QCs have also been achieved for multiple length-scale potentials with hard cores albeit more through the systematic exploration of the space of potential parameters, both in 2D [15] and 3D [16, 17].
An alternative approach to design QCs is to attempt to directly design the tendency to form a particular symmetry through directional bonding [18, 19, 20, 21, 22]. The idea is that the directionality helps to induce both local and global non-crystallographic symmetry, where the quasiperiodicity is then a consequence of the formation of that global symmetry. This approach has recently been successfully applied to design patchy particles that form icosahedral quasicrystals in simulations [21, 22]. First, it is important that the patch geometry is designed to be consistent with the desired point group symmetry; e.g. for the icosahedral QC forming systems the patches are directed along subsets of the symmetry axes of $I_{h}$. Although this encourages motifs with the desired local orientational order, it is also important that the system can form a globally ordered structure that is sufficiently stable to outcompete possible periodic crystal forms during assembly. In the icosahedral QC-forming system, this second feature was achieved by basing the patch design on the local environments in an ideal target QC (that was obtained by the cut-and-project approach), thus helping to ensure that the best way that the local environments determined by the patch geometry combine together is as a QC with the target global symmetry.
Octagonal symmetry is one of the more experimentally rarely observed quasicrystalline symmetries with only a few examples in alloys [5, 6]. Moreover, in these examples, the quasicrystalline phase is metastable, being observed to transform to a $\beta$-Mn crystal [23]. The number of examples of octagonal quasicrystal self-assembly in simulations is also modest, all in systems with isotropic interactions, both in 2D [24, 25] and in 3D [17]. Here, the goal is to increase the potential means to realize octagonal QCs by designing a system that can form a 3-dimensional octagonal QC through directional bonding.
Figure 1: (a) The ideal target octagonal quasicrystal. The structure projected down the $c$-axis is that of an Ammann-Beenker tiling with particles at the corners of the square and rhomboidal tiles. The particles are coloured by their coordination number. (b) In 3-dimensions, the square is puckered and the rhombus corresponds to a right- or left-handed helix with a pitch length of the $c$ repeat. (c) The coordination environments in the ideal quasicrystal labelled by their coordination number and their local point group symmetry. The two four-coordinate and six-coordinate environments are enantiomeric. (d) The two patchy particles. (e) The interaction matrix between the four different types of patches on the above two particles.
To apply this approach, we first need an ideal 3D octagonal QC to provide a basis for the patchy-particle designs. Our starting point is the 2D Ammann-Beenker tiling [26] that can, for example, be derived by projection of a 4D hypercubic lattice into 2D (Supplementary Section S2). The tiling consists of squares and rhombi (with internal angles of 45 and 135) with the edges of these polygons being equally likely to be oriented along eight equivalent directions. The simplest way to create a 3D octagonal QC from this tiling would be to place particles at the vertices and have a simple periodic stacking of identical Ammann-Beenker layers (as was done in Ref. [20] for a dodecagonal QC example). However, it is not feasible to realize this structure with patchy particles of well-defined radii and with bonds along the edges of the tiling because the distance across a rhomboidal tile is 0.7654 of the tile edge length and hence would lead to particle overlaps. Instead, we explored displacing particles placed at the vertices of the Ammann-Beenker tiling in $z$ so that only tile edges correspond to bonds. Our solution is presented in Fig. 1. Each bond has a component in the $z$-direction of ±$\pm c/4$ where $c$ is the periodic repeat in that direction. The rhombi become right- or left-handed helices with a pitch length of $c$, whereas the squares remain as a closed circuit of bonds, but where the pairs of diagonally-opposite particles are displaced by $c/4$ (Fig. 1(b)). We choose $c$ so that the distance between particles across the short diagonal of the projected rhombi matches the next-neighbour distance across the diagonals of the squares.
The coordination number distribution for this ideal QC is the same as for the vertices in the Ammann-Beenker tiling with local environments having from 3 to 8 neighbours. Each of the environments is a sub-environment of the 8-coordinate environment, which involves eight-equivalent bonds in a $D_{4d}$ geometry (Fig. 1(c)). The average coordination number is four.
Our basic approach to patchy-particle design is to choose the patch geometry to match the bond directions of an environment [27, 28]. For quasicrystals in particular, we have found that rather than having a different particle type for each environment (e.g. eight in the current case), the complexity of the model can be reduced by representing environments that are subsets of each other by a single patchy particle [21]. A consequence of this approach is that some of the patches would not be involved in bonds in the lower-coordinate environments, and so the particle would have an interaction energy that is lower in magnitude than its possible maximum. A concern might thus be that alternative structures that were better able to utilize all the system’s patches might be more stable.
In the current case, as all the lower-coordinate environments are subsets of the eight-coordinate, one might initially imagine a one-component system of eight-patch particles might be most suitable. However, this would lead to four unused patches per particle in the ideal quasicrystal and so have a significant potential for alternative structure formation. Instead, we initially consider a two-component system that is a mixture of five- and eight-patch particles, where it is envisaged that the eight-patch particles would be used in the 6-, 7- and 8-coordinate environments and the 5-patch particles in the 3-, 4- and 5-coordinate environments. In this case, the composition of 5- to 8-patch particles that would match the ideal quasicrystal is $(7\sqrt{2}-9)/(10-7\sqrt{2}):1\approx 8.95:1$.
Figure 2: (a) Close-up of a cut through a binary octagonal quasicrystal viewed down the 8-fold axis. Bonds are drawn between particles within $5\,\sigma_{\mathrm{LJ}}$ of the cut surface. (b) BOOD for the binary quasicrystal. (c) BOOD for a one-component quasicrystal made of 5-patch particles. BOODs were calculated with the convention that two particles are bonded if the distance between them is lower than 1.5 $\sigma_{\mathrm{LJ}}$ and the interaction energy is lower than -0.2 $\epsilon_{\mathrm{LJ}}$. BOODs are projected onto a plane perpendicular to the 8-fold symmetry axis using Lambert projection. (d) Diffraction patterns of the binary octagonal QC viewed down the 8-fold axis and a 2-fold axis. (e) Radial distribution functions for the assembled binary, P5, and P8 clusters and the ideal OQC and $C2/c$ crystal. (f) Coordination number distributions for the same systems as (e) (calculated using the same energy criterion as the BOODs in the assembled QCs).
All the patches of the 8-patch particle are equivalent by symmetry, whereas the 5-patch particle has three sets of non-equivalent patches. The specificity of the patch-patch interactions can be used to help favour the formation of the target structure. In this case, we only allow patches to interact if they would be involved in bonds in the ideal octagonal QC (the resulting interaction matrix is shown in Fig. 1(e)). For example, there are never any bonds between 6-, 7- and 8-coordinate environments in the ideal octagonal QC, so the patches on the 8-patch particle are non self-interacting.
To simulate the patchy particles, we use a modified Lennard-Jones potential where the attractive component is modulated by angular and torsional factors such that the full attractive interaction is only obtained if interacting patches point directly at each other and the particles have the correct relative orientation [27, 29, 28]. The simulations were performed with a GPU-enabled Monte Carlo algorithm [30]. The full form of the potential, the particle designs and further details about the simulations are given in the Supplementary Information.
Fig. 2(a) shows a cut through a large binary cluster grown in our simulations. From the structure it can be clearly seen that bonds are oriented along eight equivalent directions. This is confirmed by the bond-orientational order diagram (BOOD) (Fig. 2(b)) which has clear eight-fold symmetry. Similarly, the diffraction pattern (Fig. 2(d)) also exhibits eight-fold symmetry and quasiperiodic character, e.g. the ratio of the positions of the first two peaks in the $x$ direction is $\sqrt{2}$. The cluster is an octagonal quasicrystal.
In Fig. 2(f) we show the coordination number distribution for the assembled cluster. The average coordination number of our assembled octagonal QC is close to four ($\langle CN\rangle$ =3.81; see Table S6), i.e. nearly the same as the ideal QC. However, there is a relative lack of high-coordinate particles compared to the ideal Ammann-Beenker structure. Only about 1% of particles have a coordination number of six or more even though the fraction of 8-patch particles is 9%. This feature suggests that the 8-patch particles may not be actually needed for quasicrystal formation. We therefore ran simulations of a one-component system of 5-patch particles (with the same patch-patch interaction matrix (Fig. 1(e))). An octagonal quasicrystal again formed (Figs. S2, S3)) that had effectively identical structural properties to the binary system, e.g. the coordination number distribution (Fig. 2(f)) and radial distribution functions are very similar (Fig. 2(e)).
[…]
Figure 3: (a) A slab of the ideal $C2/c$ crystal viewed along the pseudo 8-fold axis. The four-coordinate environments are coloured in red and the 5-coordinate environments in green. The unit cell is shown in blue. The coordinates for the particles in the unit cell of the crystal are given in Supplementary Table S3. (b) Side views of this crystal and the ideal octagonal QC. (c) a cut of thickness 5 $\sigma_{LJ}$ through the P5 quasicrystal with the periodic direction vertical.
Examining the structure of the assembled quasicrystals, one can see features expected from the Ammann-Beeenker structure such as squares and thin rhombi, but one also sees additional motifs, such as the overlapping squares mentioned above (Fig. 2(a)). The side-view of the QC in Fig. 3(c) clearly shows greater local similarity to the crystal than the ideal octagonal QC (Fig. 3(b)). This helps to explain why the quasicrystals are able to match the coordination number of the ideal QC even though they possess considerable disorder. The additional motifs also provide a mechanism to introduce differences between the 2D quasicrystalline layers while maintaining inter-layer bonding. This is important as it has been suggested that 3D axial quasicrystals can only become stable due to their greater entropy if there is disorder in the periodic direction [34, 35]. In a stacking of identical quasicrystalline layers (as in the ideal QC of Fig. 1(a)) the random tiling entropy would only scale with the 2D area of the quasicrystal and so never be able to overcome an energy difference with respect to the most stable crystal that would scale with the volume.
One way to potentially assess the quality of the quasiperiodic order is to measure the phason strain (it is zero for a perfectly quasiperiodic structure) [16]. This analysis requires a “lifting” of the particle coordinates to 5 dimensions in a reverse of the “cut-and-project” process that can be used to generate an ideal 3D octagonal quasicrystal. However, in the current examples, this mapping is ill-defined due to the ubiquitous presence of edge dislocations in the quasicrystalline planes (the Burgers vectors for these dislocations never have any component in the periodic direction). Their presence is suggested by the bending of lines of bonds in the illustrated clusters, but can be more definitively located using the methods of Refs. [36, 37] (see Section S8a). It may well be that these dislocations provide an additional source of entropy that helps to stabilize the quasicrystals.
Finally, we also explored what happens in a one-component system of the 8-patch particles (the patches are now allowed to self-interact). The diffraction pattern and BOOD for a large cluster are shown in Fig. 4(a) and (b) and exhibit features expected of an octagonal quasicrystal. However, the situation is more complex than for the previous systems. The assembled cluster shows increased facetting (the surfaces of the quasicrystals (e.g. Fig. 3(c) and S2) are by contrast significantly rougher) and repetitive patterns. As is clear from the slices in Fig. 4(c) and (d) the cluster consists of multiple crystalline domains. Due to the coherent boundaries between the domains, a structure with an overall approximate eight-fold symmetry results. The crystalline domains are based upon the crystal illustrated in Fig. 3(a) but that has deformed to allow the additional patches that would not otherwise be involved in bonding to form bonds with some of the relatively close second neighbours (in the ideal octagonal QC the second neighbours are at a distance of only 1.2156 times the nearest neighbours (Fig. 2(e)). These extra interactions lead to the additional rings of peaks in the BOOD compared to the previous systems; other consequences are a larger average coordination number (nearly 6 (Table S6)), a shorter repeat in the periodic direction and differences in the radial distribution function (particularly associated with the second peak).
Figure 4: (a) Diffraction pattern of a 101 365 particle cluster in a one-component eight-patch system viewed down the axis of eightfold symmetry and perpendicular to it. (b) BOOD where bonds are defined when two particles are closer than 1.5 $\sigma_{\mathrm{LJ}}$ and have an energy lower than $-0.2\,\epsilon_{\mathrm{LJ}}$. (c-d) Slices through the cluster reveal the multi-domain character of the cluster. (c) is viewed down the eight-fold axis and (d) perpendicular to it.
Here, we have designed patchy particles that we have shown to be capable of assembling into an octagonal quasicrystal. An important question, though, is how might such particles be realized. Recently, DNA origami particles have been designed to assemble into increasingly complex crystals [38, 39, 40]. This has both been through DNA origami polyhedra that assemble through single strands at their vertices [38, 40] and through motifs with rigid arms that enable directional bonding [39]. Given the need for torsionally-specific interactions for our particle designs to assemble, the latter approach might be more appropriate in this case, e.g. a particle with five arms extending from a central square anti-prism. Given the impressive recent advances in protein design algorithms [41], proteins might provide another possibility [42]. For example, an approach to generate proteins that can assemble into crystals with well-defined symmetries has recently been developed [43].
Acknowledgements.
ACKNOWLEDGEMENTS: We are grateful for financial support from Agencia Estatal de Investigación through grant No. PID2020-115722GB-C21 (E.G.N), the Special Fund for Young Researchers of Keio University, provided by the Ishii Ishibashi Kikin, and by JSPS KAKENHI Grant Number JP202222673 (A.K.). We acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility.