### v2

# A patchy-particle 3-dimensional octagonal quasicrystal

Abstract
We devise an ideal 3-dimensional octagonal quasicrystal that is based upon the 2-dimensional Ammann-Beenker tiling and that is potentially suitable for realization with patchy particles. Based on an analysis of its local environments we design a binary system of 5- and 8-patch particles that in simulations assembles into a 3-dimensional octagonal quasicrystal. The local structure is subtly different from the original ideal quasicrystal possessing a narrower coordination-number distribution; in fact, the 8-patch particles are not needed and a one-component system of the 5-patch particles assembles into an essentially identical octagonal quasicrystal. We also consider a one-component system of the 8-patch particles; this assembles into a cluster with a number of crystalline domains, but which, because of the coherent boundaries between the crystallites, has approximate eight-fold order. We envisage that these systems could be realized using DNA origami or protein design.
I Introduction
Quasicrystals (QCs) have long-range order, as exemplified by the sharp peaks in their diffraction patterns, but without a periodically repeating unit cell. QCs can thus have symmetries that are not possible for periodic crystals. The first example was discovered by Shechtman for an Al/Mn alloy and had icosahedral symmetry.[1] This discovery stimulated a search for further examples and quasicrystalline alloys with decagonal,[2] dodecagonal[3, 4] and octagonal[5, 6] symmetry were quickly discovered. More recently, QCs have also been experimentally discovered in soft matter systems,[7, 8, 9] but these have been generally limited to dodecagonal symmetry. Thus, discovering new ways to form QCs and increasing the possible repertoires of structures and symmetries is of significant interest.
Theory and simulations have the potential to significantly contribute to this goal by identifying how interparticle interactions can be designed to achieve quasicrystallinity. QCs involve multiple length scales that are related by a specific irrational ratio, e.g. the golden ratio in the case of icosahedral QCs. This feature is exploited in approaches that use isotropic potentials with multiple length scales.[10, 11] For example, a theoretical framework for designing such ultra-soft potentials has been developed.[12, 13, 14] QCs have also been achieved for multiple length-scale potentials with hard cores, albeit more through the systematic exploration of the space of potential parameters, both in 2D[15] and 3D.[16, 17, 18] Inverse design methods also hold significant promise for quasicrystal discovery.[19]
An alternative approach to design QCs is to attempt to directly design the tendency to form a particular symmetry through directional bonding.[20, 21, 22, 23, 24, 25] The idea is that the directionality helps to induce both local and global non-crystallographic symmetry, where the quasiperiodicity is then a consequence of the formation of that global symmetry. This approach has recently been successfully applied to design patchy particles that form icosahedral quasicrystals in simulations.[23, 24] First, it is important that the patch geometry is designed to be consistent with the desired point group symmetry; e.g. for the icosahedral QC-forming systems the patches are directed along subsets of the symmetry axes of $I_{h}$. Although this encourages motifs with the desired local orientational order, it is also important that the system can form a globally ordered structure that is sufficiently stable to out-compete possible periodic crystal forms during assembly. In the icosahedral QC-forming system, this second feature was achieved by basing the patch design on the local environments in an ideal target QC (that was obtained by the cut-and-project approach), thus helping to ensure that the best way that the local environments determined by the patch geometry combine together is as a QC with the target global symmetry.
Octagonal symmetry is one of the more experimentally rarely observed quasicrystalline symmetries with only a few examples in alloys.[5, 6] Moreover, in these examples, the quasicrystalline phase is metastable, being observed to transform to a $\beta$-Mn-type crystal.[26] The number of examples of octagonal quasicrystal self-assembly in simulations is also modest, all in systems with isotropic interactions, both in 2D[27, 28] and in 3D.[17] Here, the goal is to increase the potential means to realize octagonal QCs by designing a system that can form a 3-dimensional octagonal QC through directional bonding.
Figure 1: The ideal target octagonal quasicrystal. (a) The structure projected down the $c$-axis is that of an Ammann-Beenker tiling with particles at the corners of the square and rhomboidal tiles. The particles are coloured by their coordination number (See Fig. 2). (b) A side-view along one of the two-fold axes. (c) In 3-dimensions, the square is puckered and the rhombus corresponds to a right- (illustrated) or left-handed helix with a pitch length of the $c$ repeat.
Figure 2: (a) The coordination environments in the ideal quasicrystal labelled by their coordination number and their local point group symmetry. The two four-coordinate and six-coordinate environments are enantiomeric. (b) The two patchy particles. (c) The interaction matrix between the four different types of patches on the above two particles.
II An ideal octagonal quasicrystal
To apply this approach, we first need an ideal 3D octagonal QC to provide a basis for the patchy-particle designs. Our starting point is the 2D Ammann-Beenker tiling [29] that can, for example, be derived by projection of a 4D hypercubic lattice into 2D (Supplementary Section S1). The tiling consists of squares and rhombi (with internal angles of 45 and 135) with the edges of these polygons being equally likely to be oriented along eight equivalent directions. The simplest way to create a 3D octagonal QC from this tiling would be to place particles at the vertices and have a simple periodic stacking of identical Ammann-Beenker layers (as was done in Ref. 22 for a dodecagonal QC example). However, it is not feasible to realize this structure with patchy particles of well-defined radii and with bonds along the edges of the tiling because the distance across the short diagonal of a rhomboidal tile is 0.7654 of the tile edge length and hence would lead to particle overlaps. Instead, we explored how to place particles at the vertices of the Ammann-Beenker tiling at different heights in $z$ so that only tile edges correspond to bonds and there are no particle overlaps. Our solution is presented in Fig. 1. Each bond has a component in the $z$-direction of $\pm c/4$ where $c$ is the periodic repeat in that direction. The rhombi become right- or left-handed helices with a pitch length of $c$, whereas the squares remain as a closed circuit of bonds, but where the pairs of diagonally-opposite particles are displaced by $c/4$ (Fig. 1(c)). We choose $c$ so that the distance between particles across the short diagonal of the projected rhombi matches the next-neighbour distance across the diagonals of the squares.
The coordination number distribution for this ideal QC is the same as for the vertices in the Ammann-Beenker tiling with local environments having from 3 to 8 neighbours. Each of the environments is a sub-environment of the 8-coordinate environment, which involves eight-equivalent bonds in a $D_{4d}$ geometry (Fig. 2(a)). The average coordination number is four.
The above proposed decoration of the square and two rhomboidal cells is clearly compatible with quasicrystal formation in systems made up of such cells, but it is interesting to ask how such a decoration constrains the ways that these cells could potentially combine. The decoration induces an edge-matching rule between the three types of cells (Fig. S2). Although this differs from the edge-matching rule of the Ammann-Beenker tiling that when combined with the appropriate vertex-matching rule ensures quasiperiodicity,[30] this is not necessarily a problem for our approach, because, like in previous patchy-particle quasicrystals,[23, 24] entropy maximization is expected to be the driving force for any quasiperiodicity observed.
This ideal QC is significantly different from previously observed 3-dimensional octagonal QCs,[5, 6, 17] as although they can also be analysed in terms of square-rhomb tilings, the decoration of the tiles with particles is much more complex, with most models of the experimental examples being related to the $\beta$-manganese structure.[31, 32] However, our ideal octagonal QC does have some structural similarities to the 3-dimensional decagonal QC observed in simulations in Ref. 17 for a multi-minimum potential in that this quasicrystal is also based on a classical quasiperiodic tiling, namely the Tübingen tiling, with particles at the vertices of the tiles and one of the tiles (the pentagon) being associated with helical configurations.
III Patchy-particle design
Our basic approach to patchy-particle design is to choose the patch geometry to match the bond directions of an environment. [33, 34] For quasicrystals in particular, we have found that rather than having a different particle type for each environment (e.g. eight in the current case), the complexity of the model can be reduced by representing environments that are subsets of each other by a single patchy particle.[23, 24] A consequence of this approach is that some of the patches would not be involved in bonds in the lower-coordinate environments, and so the particle would have an interaction energy that is lower in magnitude than its possible maximum. A concern might thus be that alternative structures that were better able to utilize all the system’s patches might be more stable.
In the current case, as all the lower-coordinate environments are subsets of the eight-coordinate environment, one might initially imagine that a one-component system of eight-patch particles might be most suitable. However, this would lead to an average of four unused patches per particle in the ideal quasicrystal and so have a significant potential for alternative structure formation. Instead, we initially consider a two-component system that is a mixture of five- (P5) and eight-patch (P8) particles, where it is envisaged that the eight-patch particles would be used in the 6-, 7- and 8-coordinate environments and the 5-patch particles in the 3-, 4- and 5-coordinate environments. In this case, the composition of 5- to 8-patch particles that would match the ideal quasicrystal is $(7\sqrt{2}-9)/(10-7\sqrt{2}):1\approx 8.95:1$ (Table S2).
Figure 3: (a) Close-up of a cut through a binary octagonal quasicrystal viewed down the 8-fold axis. Bonds are drawn between particles within $5\,\sigma_{\mathrm{LJ}}$ of the cut surface, and these particles are visualized as small green (P5) or cyan (P8) spheres. Particles further away from this cut plane are visualized as larger grey spheres. The yellow ellipse highlights an “overlapping-squares” motif (see Fig. 4(a)). (b) BOOD for the binary quasicrystal. (c) BOOD for a one-component quasicrystal made of 5-patch particles. (d) Diffraction patterns of the binary octagonal QC viewed down the 8-fold axis and a 2-fold axis. (e) Radial distribution functions for the assembled binary, P5, and P8 clusters and the ideal OQC and $C2/c$ crystal. (f) Coordination number distributions for the same systems as (e) (calculated for the assembled QCs using the same energy and distance criteria as for the BOODs).
All the patches of the 8-patch particle are equivalent by symmetry, whereas the 5-patch particle has three sets of non-equivalent patches. The specificity of the patch-patch interactions can be used to help favour the formation of the target structure. In this case, we only allow patches to interact if they would be involved in bonds in the ideal octagonal QC (the resulting interaction matrix is shown in Fig. 1(e)). For example, there are never any bonds between 6-, 7- and 8-coordinate environments in the ideal octagonal QC, so the patches on the 8-patch particle are not self-interacting. Similarly, patch 2 on the 5-patch particle would never bond to itself in the ideal quasicrystal, so we also make this patch non self-interacting.
IV Methods
IV.1 Patchy-particle potential
To simulate the patchy particles, we use a modified Lennard-Jones potential where the attractive component is modulated by angular and torsional factors such that the full attractive interaction is only obtained if interacting patches point directly at each other and the particles have the correct relative orientation.[33, 35, 34] The specific form of the interactions between the patchy particles is the same as in Refs. 35, 34. The interaction is based on a cut-and-shifted Lennard-Jones potential $V_{\mathrm{LJ}}^{\prime}$ where
$V_{\mathrm{LJ}}^{\prime}(r)=\begin{cases}V_{\mathrm{LJ}}(r)-V_{\mathrm{LJ}}(r_{\mathrm{cut}})&:r<r_{\mathrm{cut}}\\ 0&:r\geq r_{\mathrm{cut}}\end{cases}$ (1)
and
$V_{\mathrm{LJ}}(r)=4\varepsilon_{\mathrm{LJ}}\left[\left(\frac{\sigma_{\mathrm{LJ}}}{r}\right)^{12}-\left(\frac{\sigma_{\mathrm{LJ}}}{r}\right)^{6}\right].$ (2)
The patchy particle potential is described by a Lennard-Jones repulsive core and an attractive tail that is modulated by angular and torsional dependent functions:
$V_{ij}(\mathbf{r}_{ij},\mathbf{\Omega}_{i},\mathbf{\Omega}_{j})=\begin{cases}V_{\mathrm{LJ}}^{\prime}(r_{ij})&:r_{ij}<\sigma_{\mathrm{LJ}}^{\prime}\\ V_{\mathrm{LJ}}^{\prime}(r_{ij})\underset{{\mathrm{patch~pairs~}\alpha,\beta}}{\max}\left[\varepsilon_{\alpha\beta}V_{\mathrm{ang}}(\mathbf{\hat{r}}_{ij},\mathbf{\Omega}_{i},\mathbf{\Omega}_{j})V_{\mathrm{tor}}(\mathbf{\hat{r}}_{ij},\mathbf{\Omega}_{i},\mathbf{\Omega}_{j})\right]&:r_{ij}\geq\sigma_{\mathrm{LJ}}^{\prime}\end{cases},$ (3)
where $\mathbf{r}_{ij}$ is the interparticle vector, $\alpha$ and $\beta$ are patches on particles $i$ and $j$ respectively, $\mathbf{\Omega}_{i}$ is the orientation of particle $i$, $\sigma_{\mathrm{LJ}}^{\prime}$ corresponds to the distance at which $V_{\mathrm{LJ}}^{\prime}$ passes through zero and $\varepsilon_{\alpha\beta}$ is a measure of the relative strength of the interactions between patches $\alpha$ and $\beta$. We set the cutoff distance $r_{\mathrm{cut}}=2.5\,\sigma_{\mathrm{LJ}}$.
The angular modulation term $V_{\mathrm{ang}}$ is a measure of how directly the patches $\alpha$ and $\beta$ point at each other, and is given by
$V_{\mathrm{ang}}(\mathbf{\hat{r}}_{ij},\mathbf{\Omega}_{i},\mathbf{\Omega}_{j})=\exp\left(-\frac{\theta_{\alpha ij}^{2}}{2\sigma_{\mathrm{ang}}^{2}}\right)\exp\left(-\frac{\theta_{\beta ji}^{2}}{2\sigma_{\mathrm{ang}}^{2}}\right).$ (4)
$\theta_{\alpha ij}$ is the angle between the patch vector $\mathbf{\hat{P}}_{i}^{\alpha}$, representing the patch $\alpha$ on particle $i$, and $\mathbf{\hat{r}}_{ij}$. $\sigma_{\mathrm{ang}}$ is a measure of the angular width of the patch.
The torsional modulation term $V_{\mathrm{tor}}$ describes the variation in the potential as either of the particles is rotated about the interparticle vector $\mathbf{r}_{ij}$ and is given by
$V_{\mathrm{tor}}(\mathbf{\hat{r}}_{ij},\mathbf{\Omega}_{i},\mathbf{\Omega}_{j})=\exp\left(-\frac{1}{2\sigma_{\mathrm{tor}}^{2}}\left[\min\limits_{\phi^{\mathrm{offset}}_{\alpha\beta}}\left(\phi_{\alpha\beta}-\phi^{\mathrm{offset}}_{\alpha\beta}\right)\right]^{2}\right).$ (5)
where $\phi^{\mathrm{offset}}$ is the preferred value of the torsional angle $\phi$. To define the torsional angle $\phi_{\alpha\beta}$, a unique reference vector is associated with each patch. In order to capture the symmetry of an environment, more than one equivalent offset angle can be defined, in which case we find the minimum value of $\phi-\phi_{\mathrm{offset}}$ across the set of equivalent offset angles. The torsional interaction prevents free rotation about a patchy bond and encourages particles to bind with the correct relative orientation, thus facilitating the propagation of long-range orientational order.
As in previous work we choose $\sigma_{\mathrm{tor}}=2\,\sigma_{\mathrm{ang}}$. We also choose $\sigma_{\mathrm{ang}}$ to be sufficiently narrow to give strong directional bonding, but not too narrow as to hinder the kinetics. We use both $\sigma_{\mathrm{ang}}=0.2$ and 0.3 radians. $\varepsilon_{\alpha\beta}$ is either 0 or 1, matching the interaction matrix in Fig. 1(e); as this simple choice was successful in achieving quasicrystal assembly, there was no attempt to further optimize the strengths of the interacting patches. We use $\sigma_{\mathrm{LJ}}$ (the distance at which the Lennard–Jones potential is zero) as our unit of length, and the Lennard–Jones well depth $\varepsilon_{\mathrm{LJ}}$ as our unit of energy. Temperatures are given in reduced form, i.e. $T^{*}=k_{B}T/\varepsilon_{\mathrm{LJ}}$.
The details of the patchy particle designs for the binary system proposed in Section III are given in Table S4. The symmetry axis of the 8-patch particle is used as the reference vector for the torsional component of the potential. An offset angle of 180 ensures that the preferred relative geometry is with the eight-fold axes of the particles alligned. The patches of the 5-patch particle are a subset of those of the 8-patch particle.
IV.2 Simulations
The simulations were performed with a GPU-enabled Monte Carlo algorithm.[36] Our basic simulation protocol is similar to our previous work on 3-dimensional patchy-particle quasicrystals.[23, 22, 24] Namely, assembly is started from a low-density fluid (with 20 000 particles), choosing a temperature that is just sufficiently low enough to allow nucleation of the ordered phase (we generally want to avoid multiple nucleation events). The composition of the fluid in the binary system was chosen to match that for the ideal target quasicrystal. Once we have grown a moderately-sized cluster (typically of order 10 000 particles) we then place this cluster in a larger box (and with a larger reservoir of particles in the fluid phase) and continue growth. We may increase the temperature slightly in this second phase to avoid any further nucleation events. For the results reported in Section V.1 and V.2, we typically ran the simulations until a cluster with close to 100 000 particles was obtained. More details of these simulations are given in Table 1.
For the simulations reported in Section V.3, which test the effects of varying potential and design parameters, instead the clusters were usually grown just until they were sufficiently large for QC formation to be confidently determined from the diffraction patterns and BOODs; this corresponded to at least 10 000 particles.
[…]
max.
system $\sigma_{\mathrm{ang}}$ seed? $T_{\mathrm{init}}$ $T_{\mathrm{growth}}$ cluster size $\langle CN\rangle$
binary 0.2 none 0.078 0.078 80 911 3.81
binary 0.2 ideal OQC 0.078 0.079 79 800 3.76
P5 0.3 none 0.0910 0.0915 82 330 4.03
P5 0.3 $C2/c$ crystal 0.0910 0.0920 89 839 4.03
P8 0.3 none 0.11 0.12 101 365 5.93
[…]
One useful way to characterize the order in the systems is using a bond orientational order diagram (BOOD). The BOOD is a plot of the first coordination shell of each particle on a unit sphere, which is subsequently projected onto a plane using an area-preserving Lambert projection. We choose to perform the projection onto a plane plane perpendicular to the 8-fold symmetry axis.
V Results
V.1 Quasicrystalline binary and P5 systems
Fig. 3(a) shows a cut through a large binary cluster of P5 and P8 particles grown in our simulations. From the structure it can be clearly seen that bonds are oriented along eight equivalent directions (images of the complete cluster are shown in Fig. S5) and no periodicity is apparent. The former feature is further confirmed by the bond-orientational order diagram (BOOD) (Fig. 3(b)) which has clear eight-fold symmetry. Similarly, the diffraction pattern (Fig. 3(d)) also exhibits eight-fold symmetry and quasiperiodic character, e.g. the ratio of the positions of the first two peaks in the $x$ direction is $\sqrt{2}$. The cluster is an octagonal quasicrystal.
In Fig. 3(f) we show the coordination number distribution for the assembled cluster. The average coordination number of our assembled octagonal QC is close to four (Table 1), i.e. nearly the same as the ideal QC. However, there is a relative lack of high-coordinate particles compared to the ideal Ammann-Beenker structure. Only about 1% of particles have a coordination number of six or more even though the fraction of 8-patch particles is 9%. This feature suggests that the 8-patch particles may not be actually needed for quasicrystal formation. We therefore ran simulations of a one-component system of 5-patch particles (with the same patch-patch interaction matrix (Fig. 1(e))). An octagonal quasicrystal again formed (Figs. S5 and S7)) that had very similar structural properties to the binary system, e.g. the coordination number distribution (Fig. 3(f)) and radial distribution functions are very similar (Fig. 3(e)).
[…]
Due to the larger coordination number, the crystal is significantly lower in energy than the octagonal quasicrystals, raising the possibility that the observed quasicrystals might just be kinetic products. In the absence of spontaneous transitions between crystals and quasicrystals,[37, 38] ascertaining whether quasicrystals are thermodynamically stable or just metastable is a difficult challenge [39] and one we leave to future work. Interestingly, when a growth simulation is started from a 279-particle crystalline seed the cluster that results has octagonal QC character (Fig. S7 and S11) and, aside from the seed, a local structure that is essentially identical to the unseeded P5 system. The crystal thus is able to template the growth of the quasicrystal. The preference for the quasicrystal is presumably because it has a higher growth rate, which in turn is because the greater configurational entropy inherent to the quasicrystal means there are more ways that adsorbing particles can be added that are consistent with the growth of the quasicrystalline phase.
Similarly, we also considered the growth of the binary system around a seed of the ideal octagonal quasicrystal. Unsurprisingly, this also resulted in an octagonal quasicrystal, again with negligible difference from the quasicrystal produced by homogeneous nucleation in the binary system (Fig. S7 and S11).
Figure 4: (a) A slab of the ideal $C2/c$ crystal viewed along the pseudo 8-fold axis. The 4-coordinate environments are coloured in red and the 5-coordinate environments in green. A primitive unit cell is shown in blue (the six particles in this cell are shown as spheres). An “overlapping squares” motif is highlighted. The coordinates for the particles in the conventional unit cell of the crystal are given in Supplementary Table S3. The diffraction patterns for this crystal are shown in Fig. S9. (b) Side views of this crystal and the ideal octagonal QC. (Further side views of the crystal are given in Fig. S4.) (c) A cut of thickness 5 $\sigma_{\mathrm{LJ}}$ through the P5 quasicrystal with the periodic direction vertical. Note that the features that look like bow ties in (b) and (c) are side views of the square motifs (Fig. 1(c)).
Examining the structure of the assembled quasicrystals, one can see features expected from the Ammann-Beeenker structure such as squares and thin rhombi, but one also sees additional motifs, such as the overlapping squares mentioned above (Fig. 3(a) and Fig. 4(a)). The side-view of the QC in Fig. 4(c) (see Fig. S5 for a similar cut of the binary system) clearly shows that the order in the periodic direction has greater local similarity to the crystal than the ideal octagonal QC (Fig. 4(b)). This helps to explain why the quasicrystals are able to match the coordination number of the ideal QC even though they possess considerable disorder. The additional motifs also provide a mechanism to introduce differences between the 2D quasicrystalline layers while maintaining inter-layer bonding. This is important as it has been suggested that 3D axial quasicrystals can only become stable due to their greater entropy if there is disorder in the periodic direction.[40, 41] In a stacking of identical quasicrystalline layers (as in the ideal QC of Fig. 1(a)) the random tiling entropy would only scale with the 2D area of the quasicrystal and so never be able to overcome an energy difference with respect to the most stable crystal that would scale with the volume.
Figure 5: (a) A slab of dimensions of $20\times 20\times 4$ (in units of $\sigma_{\mathrm{LJ}}$) from the binary quasicrystal with the particles represented as points (P5: green; P8: cyan). If viewed at a low angle, it becomes easier to see the lines of points and hence to spot the dislocations. (b) Inverse Fourier-transformed image of two equal and opposite diffraction spots in the first intense ring of the Fourier transform of the above slab. Edge dislocations are indicated by a ‘T’. Circuits around the dislocations are drawn in blue. If these are followed the number of lines on one side of the dislocation will be different from the other. (Plots for the three other pairs of equivalent diffraction spots are shown in Fig. S10 and enable further dislocations to be identified.)
A detailed examination of the configurations of the assembled quasicrystals for the binary and one-component 5-patch systems also reveal features (e.g. lines of particles that terminate and the bending of lines of particles near to the termination point) that are suggestive of edge dislocations in the quasicrystalline planes. However, such features are harder to interpret with confidence for quasicrystals than for periodic crystals just by visual inspection. To obtain a more definitive identification of the presence of dislocations, we follow the approach introduced in Refs. 42, 43. In this approach, two peaks at $\mathbf{q}$ and $\mathbf{q}$ in the diffraction pattern corresponding to planes of the appropriate orientation are selected and then an inverse Fourier transform performed with the rest of the pattern masked. If no dislocations are present, a series of parallel lines will be obtained. However, a dislocation is revealed by the termination of one of these lines. More specifically, if a circuit is drawn around the dislocation, the number of lines on either side of the dislocation will differ by 1.
Fig. 5(b) shows such an image for one of the four equivalent directions in the binary octagonal quasicrystal generated from a thin (along the periodic direction) slab. Edge dislocations can be clearly located in the image. By contrast, we never observed any dislocations associated with the periodic direction. This approach also confirms the presence of edge dislocations in the one-component 5-patch octagonal QC.
An edge dislocation is a line defect in a 3D periodic crystal, but a point defect in a 2D periodic crystal. In the limit that the structure within each successive quasicrystalline plane is uncorrelated, the dislocations would just be point defects specific to that plane. In our quasicrystals there are correlations between successive planes in order to facilitate interplane bonding, but, as noted above, the range of this order is relatively short. This disorder means that one can only follow the path of an edge dislocation in the quasicrystal over a relatively short distance, rather than through the whole structure until it exits the surface or forms a complete loop as is typical in periodic crystals.
It is interesting to ask why these dislocations seem so common in the current octagonal systems, but are completely absent in the patchy-particle icosahedral quasicrystals that we have previously studied.[23, 24] It may reflect that the icosahedral QC is quasiperiodic in all three dimensions, whereas the competition between the different ways of propagating the order in the periodic direction for the octagonal QCs likely play a key role. In particular, domains of different type of order in the periodic direction are visible in the vertical cuts through the assembled quasicrystals (Fig. 4(c) and Fig. S5) that correspond to the different potential orientations of the crystal-like ordering (Fig. S4). Where these meet, edge dislocations may arise. Given their ubiquity, one would presume that in the current octagonal systems the free-energetic cost of the dislocations must be relatively low. This may well mean that they are thermodynamically stable, rather than just being a consequence of the kinetics of growth, and so, contribute to the entropic stabilization of the QC phase.
One way to potentially assess the quality of the quasiperiodic order is to measure the phason strain (it is zero for a perfectly quasiperiodic structure).[16] This analysis requires a “lifting” of the particle coordinates to 5 dimensions in a reverse of the “cut-and-project” process that can be used to generate an ideal 3D octagonal quasicrystal. However, in the current examples, this mapping is ill-defined due to the ubiquitous presence of edge dislocations in the quasicrystalline planes.
Figure 6: (a-b) Diffraction pattern of a 101 365 particle cluster in a one-component eight-patch system viewed (a) down the axis of approximate eight-fold symmetry and (b) perpendicular to it. (c) BOOD where bonds are defined when two particles are closer than 1.5 $\sigma_{\mathrm{LJ}}$ and have an energy lower than $0.2\,\varepsilon_{\mathrm{LJ}}$. (d) For each spot in the BOOD, the percentage deviation from the mean (averaged over symmetry-equivalent spots) value of the integral under the peak for each of the assembled QCs. We divide the 32 spots in the P8 BOOD into two types; namely, the more intense spots (P8) that are analogous to those that appear in the P5 and binary QCs and the less intense spots that appear due to the additional structural distortion present in the P8 system (P8 new spots). (e-f) Slices through the cluster reveal the multi-domain character of the cluster. (e) is viewed down the approximate eight-fold axis and (f) perpendicular to it.
V.2 P8 system
We also explored what happens in a one-component system of the 8-patch particles (the patches are now allowed to self-interact). The diffraction pattern and BOOD for a large cluster are shown in Fig. 6 and exhibit features similar to that of an octagonal quasicrystal. However, the situation is more complex than for the previous systems. The assembled cluster shows increased facetting (the surfaces of the quasicrystals (e.g. Fig. 4(c) and S5) are by contrast significantly rougher) and repetitive patterns that suggest local periodicity (Fig. 6 and S5). As is clear from the slices in Fig. 6(e) and (f) the cluster consists of multiple crystalline domains. Due to the coherent boundaries between the domains, a structure with an overall approximate eight-fold symmetry results. The crystalline domains are based upon the crystal illustrated in Fig. 4(a) but that has deformed to allow the additional patches that would not otherwise be involved in bonding to form bonds with some of the relatively close second neighbours (in the ideal octagonal QC the second neighbours are at a distance of only 1.2156 times the nearest neighbours (Fig. 3(e)). These extra interactions lead to the additional rings of eight smaller peaks in the BOOD that are not present in the previous systems and that occur closer to the symmetry axis. Other consequences of this change in local structure are a larger average coordination number (nearly 6 (Table 1)), a shorter repeat in the periodic direction (Fig. S7) and differences in the radial distribution function. In particular, there is now no longer a peak close to $1.36\,\sigma_{\mathrm{LJ}}$ because the distortion causes some of the second neighbours to be incorporated into the first peak and the others to be pushed to larger separations (Fig. 3(e)).
If one looks closely at the set of smaller peaks in the BOOD associated with the new patchy bonds, one can notice that the peaks are not all of the same intensity, i.e. the system does not have perfect 8-fold order. Integrating the area under the peaks confirms this conclusion and provides a more quantitative measure of the deviation from perfect symmetry (Fig. 6(d)). The integrated area associated with the weakest peaks in this set are nearly half that for the strongest. Systematic intensity deviations are also present in the other sets of peaks but these are significantly smaller in magnitude. By contrast, the intensity of the BOOD peaks for the binary and 5-patch systems are essentially identical aside from a small amount of statistical noise, further confirming their eight-fold symmetry. One interesting question is whether close to perfect 8-fold order might be restored for larger P8 clusters if the cluster size is much larger than the domain size of the crystallites. However, the large simulations that would be required are beyond the scope of this work.
V.3 Effects of potential and particle parameters
An important question, particular when considering experimental realization, is how sensitive is the observed behaviour to the potential and particle design parameters. Here, we begin to address this issue by considering some of the effects of patch specificity, patch width, the presence of torsions and the patch positions.
In all the simulations of the one-component 5-patch system discussed so far, the specificity of the patch-patch interactions has been that shown in Fig. 1(e); i.e. patch 2 cannot interact with itself. We also considered a simpler system where all the patches of the P5 particles could interact with any other; such a system might be easier to realize experimentally. We found that removal of this constraint had little effect on the assembly and an octagonal QC again resulted.
We also tested how sensitive the propensity of these systems to form an octagonal quasicrystal is on patch width. The effects of the patch width on the self-assembly of a variety of classes of ordered structures, be they symmetric complexes,[33, 35] crystals [44] or quasicrystals,[23] have been previously characterized. If a structure can exactly match the patch geometry, the general expectation is that the structure will remain stable as the patch width decreases but the kinetics of assembly slows down, whereas as the patch width increases there will come a point at which the potential is insufficiently directional to favour the target structure. To locate the latter limit we considered increments of 0.1 in $\sigma_{\mathrm{ang}}$. In the binary and the standard P5 system $\sigma_{\mathrm{ang}}=0.4$ radians was the last value at which structures with eight-fold order assembled, albeit with noticeably greater disorder. Interestingly, for the P5 variant with lower specificity eight-fold order in $x$ and $y$ was lost at $\sigma_{\mathrm{ang}}=0.4$ but layering was still seen in $z$; the local structure was quite similar to the octagonal QC but long-range orientational order was absent. These results are consistent with the idea that systems with greater interaction specificity can tolerate greater angular flexibility in their interactions.
As noted in Section V.2, the one-component P8 systems increase their coordination number by adopting a structure in which the patches are no longer able to perfectly point at each other. The energetic penalty for this deviation from perfect alignment will increase as the patch width decreases. Indeed, at $\sigma_{\mathrm{ang}}=0.2$ the system instead forms planar sheets of puckered squares. These are able to achieve a coordination number of 4, whilst having significant entropy associated with the fluctuations of the sheets. The latter is the likely reason that these 2-dimensional assemblies are observed rather than one of the 3-dimensional structures with similar coordination number. Note that the patch geometry of the P5 particles is not compatible with the pattern of bonding in these sheets and so this alternative assembly is only relevant to the P8 systems.
In all the simulations described so far, the patchy-particle potential has involved a torsional component (Eq. 5) that favours interacting particles to adopt relative orientations that match the target structure, namely with the reference vectors for the interacting patches anti-parallel (Table S4). For some of the potential approaches to realizing patchy particles, it may be hard to generate a torsional component to the potential. Therefore, we also considered the assembly behaviour with the torsional component absent. However, unlike some of the systems that form icosahedral QCs,[23, 24] it was never feasible to grow any octagonal QCs without the torsional interactions present. Instead, disordered configurations always resulted. The torsional component of the potential must sufficiently disfavour otherwise feasible alternative configurations that octagonal QC formation results.
Finally, we also considered the effects of deviations in the patch geometry. The specific deviation was an additional rotation $\alpha$ between the upward- and downward-facing patches. For the P8 particle, such a distortion would reduce its symmetry from $D_{4d}$ to $D_{4}$ and for the P5 particles from $C_{s}$ to $C_{1}$, thus making both particles chiral. We note that patchy particles made from proteins or DNA cannot possess any mirror symmetry. We specifically considered the effects of this distortion on systems of P5 particles. For $\alpha$=5 the system is still able to assemble into an octagonal QC (Fig. 7(a)). Interestingly, the BOOD retains its $D_{4d}$ symmetry albeit with slightly wider peaks than for the undistorted P5 system. Although the particles are chiral, they adapt themselves to form a structure that is on average achiral. We note that for the ideal octagonal QC based on the Ammann-Beenker tiling it is clear that this structure cannot coherently distort to accommodate this change in particle geometry.
By contrast at $\alpha$=10 the assembled cluster has begun to lose its global octagonal order. Although the local structure of the cluster is similar, the overall structure seems to have split into domains that are imperfectly aligned with each other (Fig. S12); this is reflected in the BOOD where the eight peaks are not only much wider but seem to consist of multiple sub-peaks that presumably originate from these different domains (Fig. 7(b)). These changes are likely a reflection of the greater energetic cost associated with the mis-alignment between the patch geometry and the structure of the octagonal QC.
Figure 7: BOOD and diffraction patterns of clusters of about 75 000 particles assembled from one-component chiral P5 particles (in which all patches are allowed to interact among themselves) obtained by (a) a 5 and (b) a 10 rotation between the upward- and downward- facing patches.
VI Conclusions
Here, we have designed patchy particles that we have shown to be capable of assembling into a 3-dimensional octagonal quasicrystal. Although the patchy-particle designs stemmed from an ideal 3-dimensional octagonal quasicrystal based on the Ammann-Beenker tiling, the structures of the assembled octagonal quasicrystals were subtly different, having a narrower coordination number distribution and a local structure, particularly in the periodic direction, that more resembled an alternative crystal structure. The former meant that the 8-patch particles that were part of the original particle design were not necessary for quasicrystal formation and, so, quasicrystals could be formed from a one-component system of 5-patch particles. It is noteworthy that formation of the $C2/c$ crystal was never observed in any of the simulations of these two systems, even though it will be significantly lower in free energy than the quasicrystal at sufficiently low temperature. Furthermore, although local motifs similar to that in the crystal are observed, they only extend over short length scales in the quasicrystalline planes.
The current paper is part of a programme of research seeking to understand to what extent particles with directional interactions can be used to direct the formation of 3-dimensional quasicrystals. The longer term goal is that this would then facilitate the expansion of both the types of systems for which quasicrystals can be achieved experimentally and the range of symmetries they exhibit. The current results have thus increased the symmetries of 3D patchy-particle QCs beyond just the dodecagonal [22] and icosahedral [23, 24] examples that have been developed previously to now include octagonal QCs. This is particularly significant as the number of octagonal quasicrystals observed experimentally is relatively limited and it is not clear if any are actually thermodynamically stable. Furthermore, the structure of the current octagonal quasicrystals are quite different from those previously observed in experiments[5, 6] or simulations.[17]
The driver of the quasiperiodicity is likely to be entropy maximization, like in random tiling models of quasicrystals,[45] but where the sources of entropy are more extensive. Whether the octagonal QCs discovered here are thermodynamically stable (i.e. lowest in free energy over a certain temperature range) or just a kinetic product is also an important, but challenging, question. Aspects of our results offer encouragement for the idea that the current octagonal QCs may be thermodynamically stable. Firstly, the quasicrystals exhibit significant disorder in the periodic direction, i.e. they are not identical repeats of a set of quasicrystalline planes. The above is thought to be a necessary prerequisite for the entropy of axial QCs to scale with the number of particles and hence for the QCs to be able to compete thermodynamically with an energetically more stable crystal.[40, 41] Secondly, we never see crystal formation in the binary or P5 systems. Thirdly, we observed growth of an octagonal QC from a crystal seed. Although the latter two could be the results of purely kinetic effects, such features become more likely as the thermodynamic stability of the QC with respect to the crystal increases. In particular, if the preferential growth of an octagonal QC on the crystal persists up to the melting point, then the QC is thermodynamically more stable.
The likely best way to address this question of thermodynamic stability would be by “direct coexistence” simulations [46, 47] to determine the melting point of the quasicrystal compared to potential competing crystals. This approach requires the use of a “slab” geometry where the slab crosses the periodic boundaries of the cell in two of the three dimensions. One potential issue is that the formal incompatibility of these boundary conditions with the symmetry of the quasicrystal is likely to lead to defects being present in the quasicrystal,[24] but for a sufficiently large box the consequent effects on the melting point may be relatively minor.
Another important question is how might particles analogous to those studied here be realized. Recently, DNA origami particles have been designed to assemble into increasingly complex crystals.[48, 49, 50, 51] This has both been through DNA origami polyhedra that assemble through single strands at their vertices [48, 50, 51] and through motifs with rigid arms that enable directional bonding.[49] Given the need for torsionally-specific interactions for our particle designs to assemble, the latter approach might be more appropriate in this case, e.g. a particle with five arms extending from a central square anti-prism. Given the impressive recent advances in protein design algorithms,[52] proteins might provide another possibility.[53] For example, an approach to generate proteins that can assemble into crystals with well-defined symmetries through the hierarchical assembly of symmetric clusters has recently been developed.[54] Using such an approach analogues of the higher-symmetry P8 particles are likely to be easier to realize, however, there have been also significant recent advances in assembling complexes with programmed symmetry breaking,[55, 56] as well as more arbitrary multi-component complexes using standardized protein building blocks.[57] A more detailed discussion of potential realization strategies is provided in Section S7.
Supplementary Material
The supplementary material provide further details on (i) the ideal QC, including its higher-dimensional description, the local environments and its crystalline approximants, (ii) the patchy-particle designs, (iii) the alternative monoclinic crystal, (iv) additional structural characterization of the assembled QCs and (v) potential approaches to realize experimental analogues of the patchy particle.
Acknowledgements.
We are grateful for financial support from Agencia Estatal de Investigación through grants No. PID2020-115722GB-C21 and PID2023-151751NB-I00 funded by MCIN/AEI/10.13039/501100011033 (E.G.N), the Special Fund for Young Researchers of Keio University, provided by the Ishii Ishibashi Kikin, and by JSPS KAKENHI Grant Number JP202222673 (A.K.). We acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility.
Author Declarations
Conflict of Interest
The authors have no conflicts to disclose.
Data Availability
The code used to perform the Monte Carlo simulations is available at https://github.com/evanoya/MC_GPU. Data associated with the project is available at the Oxford University Research Archive (ORA) at https://dx.doi.org/10.5287/ora-eorewogw8.